OpenStudy (anonymous):

Given 2x – 3y = 6 find the missing value in each ordered pair. (0, ), ( ,0), (6, ), ( , -4 )

OpenStudy (kainui):

Ahh, show me your best guess and we'll go from there. Remember that in an ordered pair the first number is the x-coordinate and the second one is the y-coordinate. Since they're linked by an equation, if you know one, you already have a way to find the other.

OpenStudy (anonymous):

I honesty have no idea what to do. I thought about putting the 0 in for the xand making it 2(0)-3y=6 but i dont know what to do after that

OpenStudy (tkhunny):

Obviously, you DO have an idea. The substitution was fine. Now, solve for y. What else would you do?

OpenStudy (anonymous):

thats where im stuck

OpenStudy (tkhunny):

Okay, solve this for w 2w = 4

OpenStudy (anonymous):

you would divide 2w by 2 and 4 by to so w=2

OpenStudy (tkhunny):

Okay, solve this for s. s - 4 = 3

OpenStudy (anonymous):

add 4 to the -4 and add 4 to the 3. s=7

OpenStudy (tkhunny):

There you go. You have all the skills you need to solve this expression for y. 2(0)-3y=6 Go!

OpenStudy (anonymous):

well i think the 0 goes away because it doesn't make a difference when your subtracting... 3y=6 so i just divide 3y by 3 and 6 by 3 so y=2

OpenStudy (tkhunny):

So close!! 2(0) - 3y = 6 0 - 3y = 6 -3y = 6 y = -2 You just lost a sign in there. Sweet! Do the next value.

OpenStudy (anonymous):

so i think it would make it 2x-3(0)=6 2x-0=6 2x=6 then 2x/2 amd 6/2 x=3

OpenStudy (tkhunny):

Yes!